the prod. of 2 consec. pos. intergers is 29 more than the sum. find the intergers. i havbe 5,6 is this right?
Answer by scott8148(5891)

5 times 6 is 30 ___ 5 plus 6 is 11 ___ difference is 19, not 29

x (x + 1) = x + x + 1 + 29

x^2 + x = 2x + 30

x^2 - x - 30 = 0

(x + 5) (x - 6) = 0

x = -5 and x = 6

the integers are 6 and 7 (you got 1)